英語 (English)

基盤英語 (Basic English) (医 ((保 C)1 年))

バロッグスザンネ · 非常勤講師 / 全学共通教育センター 1 単位前期水 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】 English has become the lingua franca in today’s global society. As globalization brings the world closer together, the need for global citizens to be able to communicate clearly in English is becoming more urgent. The aim of this course is to give students the opportunity to practice speaking English with an emphasis on correct pronunciation, rhythm and intonation. Listening to authentic texts by native speakers (through DVD and other media) will be an important part of the course. Practice in the four skills reading, writing, listening and speaking will be provided during class and through homework.

【授業の概要】 Students will be expected to take part in a number of role-plays and memorize short skits. Emphasis will be placed on pronunciation with natural rhythm and intonation. The following schedule is tentative and shows only general topics for each class. It may be modified depending on students’ level and interest.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. Students will gain confidence in their understanding of English pronunciation (linking, assimilation, rhythm and intonation).

2. Students will improve their listening comprehension. 【授業の計画】

1. Introduction: self introductions

2. Listening activity and role play: Harry meets Ron

3. Listening to Narnia 1: Lucy meets Mr. Tumnus

4. Listening to Narnia 2: Lucy meets Mr. Tumnus (Part 2)

5. Listening to Narnia 3: Tea at Mr. Tumnus’House

6. Listening to Narnia 4: Back through the wardrobe

7. Listening to Narnia 5: Turkish delight

8. Listening to Narnia 6: Turkish delight (Part 2) and Lucy finds Edmund

9. Listening Test 1 10. Listening to Narnia 7: A Christmas Present

11. Listening to Narnia 8: Edmund at the Witch’s House

12. Listening to Narnia 9: Aslan’s Camp

13. Listening to Narnia 10: The Witch comes to Aslan’s Camp

14. Presentations

15. Listening Test 2

16. Consolidation

【教科書】 Handouts will be prepared by the instructor and no textbook is required. 【参考書等】 [参考資料] 【成績評価の方法】 Students will be assessed on class participation, homework, role plays, skits and written assignments. Tentative coursework evaluation: Attendance: 30%, Listening Tests: 40%, Written Assignment: 15%, Presentation 15%

【再試験の有無】 Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220836 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ バロッグ .

English

Basic English (医 ((保 C)1 年))

Susanne Balogh · Part-time Lecturer / Center for General Education 1 unit 前期水 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ English has become the lingua franca in today’s global society. As globalization brings the world closer together, the need for global citizens to be able to communicate clearly in English is becoming more urgent. The aim of this course is to give students the opportunity to practice speaking English with an emphasis on correct pronunciation, rhythm and intonation. Listening to authentic texts by native speakers (through DVD and other media) will be an important part of the course. Practice in the four skills reading, writing, listening and speaking will be provided during class and through homework.

Outline⟩ Students will be expected to take part in a number of role-plays and memorize short skits. Emphasis will be placed on pronunciation with natural rhythm and intonation. The following schedule is tentative and shows only general topics for each class. It may be modified depending on students’ level and interest.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. Students will gain confidence in their understanding of English pronunciation (linking, assimilation, rhythm and intonation).

2. Students will improve their listening comprehension.

Schedule⟩

1. Introduction: self introductions

2. Listening activity and role play: Harry meets Ron

3. Listening to Narnia 1: Lucy meets Mr. Tumnus

4. Listening to Narnia 2: Lucy meets Mr. Tumnus (Part 2)

5. Listening to Narnia 3: Tea at Mr. Tumnus’House

6. Listening to Narnia 4: Back through the wardrobe

7. Listening to Narnia 5: Turkish delight

8. Listening to Narnia 6: Turkish delight (Part 2) and Lucy finds Edmund

9. Listening Test 1 10. Listening to Narnia 7: A Christmas Present

11. Listening to Narnia 8: Edmund at the Witch’s House

12. Listening to Narnia 9: Aslan’s Camp

13. Listening to Narnia 10: The Witch comes to Aslan’s Camp

14. Presentations

15. Listening Test 2

16. Consolidation

Textbook⟩ Handouts will be prepared by the instructor and no textbook is required. Reference⟩ [参考資料] Evaluation Criteria⟩ Students will be assessed on class participation, homework, role plays, skits and written assignments. Tentative coursework evaluation: Attendance: 30%, Listening Tests: 40%, Written Assignment: 15%, Presentation 15%

Re-evaluation⟩ Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220836 Contact (Office-Hour, Room, E-mail)⟩

⇒ Balogh .